OpenStudy (anonymous):

patrick has 15 coins, some dimes and some quarters. if their value is $2.25, how many of each does he have???????

OpenStudy (amistre64):

some amount of dimes + some amount of quarters = 15 coins d + q = 15 ; means that d = 15-q some number of d(.10) + some number of q(.25) = $2.25

OpenStudy (anonymous):

is it 5q and and 10d

OpenStudy (amistre64):

(15-q)(.10) + q(.25) = 2.25 1.50 -.10q +.25q = 2.25

OpenStudy (amistre64):

.15q = .75 q = .75/.15 = 15/3 = 5 5qs

OpenStudy (amistre64):

which then of course means 10 dimes

OpenStudy (anonymous):

Let x and y be the number of dimes and quarters respectively, then \(x+y=15\) and \(0.1x+0.25y=2.25\). From equation (1) we have \(x=15-y\), so \(0.1(15-y)+0.25y=2.25 \implies 1.5+0.15y=2.25 \implies y=5\). Therefore \(x=10\). So patrick has 10 dimes and 5 quarters.

OpenStudy (anonymous):

so i got it right ty {|: